OpenStudy (anonymous):

how do you find the square root of a fraction? I really can't figure it out, can someone help me please?

OpenStudy (anonymous):

take the square root of the top and bottom

OpenStudy (anonymous):

the denominator is a decimal, and i still can't figure it out..

OpenStudy (anonymous):

what is the question?

OpenStudy (jamesj):

Remember first that \[ \sqrt{ab} = \sqrt{a}\sqrt{b} \] From that it follows that \[ \sqrt{\frac{a}{b}} = \sqrt{a \frac{1}{b}} = \sqrt{a}\sqrt{\frac{1}{b}} = \frac{\sqrt{a}}{\sqrt{b}} \] For example, \[ \sqrt{\frac{4}{9}} = \frac{\sqrt{4}}{\sqrt{9}} = \frac{2}{3} \]

OpenStudy (anonymous):

the question is asking : To the nearest hundredth of a second, how long would it take for an object dropped from a height of 50 meters to land on the Moon? You're suppose to find the square root of 50/0.82

OpenStudy (jamesj):

Just use a calculator then: ccalculate 50/0.82, then take the square root of that.

OpenStudy (anonymous):

thank you sooooo much!!

OpenStudy (anonymous):

right

OpenStudy (anonymous):

love the drawing